# Parameter-Efficient Fine-Tuning in Large Models: A Survey of Methodologies

### v2

@@ V Future Directions
    PEFT methods for multi-objective tasks: In the real world, downstream tasks that LLMs need to adapt to may involve multiple objectives. For example, in paper [234], the authors emphasize that for the Program Repair task, it is necessary to consider two objectives simultaneously: 1) adapt the LLM parameters to the syntactic nuances of the task of code transformation, and 2) specifically fine-tune the LLM with respect to the logical reason behind the code change in the training data. Since most existing PEFT methods are not designed for multi-objective scenarios, exploring and developing PEFT methods suitable for multi-objective tasks is an important direction for future research.
@@
    PEFT methods in multimodal learning: The majority of existing PEFT methods have been architected with a focus on single-modality LLMs. However, as the field of LLMs evolves, the emergence of multimodal LLMs has garnered increasing interest within the research community. As observed in the experimental findings of [235], fine-tuning the connector layers, a critical component in multimodal large language models, does not always yield better results; its effectiveness depends on specific circumstances. This suggests that PEFT methods for multimodal large language models warrant further exploration and research.
@@
    Automated design of adapter modules: The adapter module is characterized by the strategic intercalation of diminutive adapter layers within the Transformer architecture’s blocks. Typically, an adapter module initiates with a down-projection matrix, proceeds through a nonlinear activation function, and culminates with an up-projection matrix. The bottleneck dimensionality within these projection matrices is a hyperparameter that often necessitates manual tuning and is potentially contingent upon the specific task requirements [117, 115]. Consequently, there is a critical imperative to devise algorithms capable of dynamically adjusting these hyperparameters in accordance with task-specific information, thereby optimizing the adapter’s efficacy across diverse applications.
@@
    Heuristic search strategies for hybrid PEFT methods: The effectiveness of different PEFT techniques can vary greatly depending on the task at hand. Consequently, many researchers are working to integrate the benefits of various PEFT strategies. However, most existing hybrid methods require the selection of PEFT methods and combination modes in advance and cannot adapt to task differences autonomously. For example, in paper [168], the authors, under a predefined design space, conduct numerous experiments to determine an ideal hybrid strategy. However, the optimal hybrid strategy may not be included within this artificially predefined design space. Therefore, introducing heuristic search strategies to find the best hybrid strategy is a promising direction for future research.
@@
    Continual learning for PEFT methods: To enhance the alignment of pre-trained large models with discrete downstream applications, the prevailing strategy entails leveraging a modestly-sized annotated dataset specific to the target task for the purpose of Supervised Fine-Tuning. In the architectural phase of PEFT methods, there exists a paucity of focus on the preservation or augmentation of the pre-trained model’s capacity to recall and effectively leverage its embedded knowledge corpus. This oversight may prove detrimental in scenarios where downstream tasks are characterized by frequent data revisions or are subject to swift environmental fluctuations. Consequently, the integration of Continual Learning principles within the framework of PEFT methods represents a significant avenue for scholarly exploration and advancement.
@@
    Differential privacy for PEFT methods: Different downstream tasks often involve varying levels of sensitve and personal data, which further emphasizes the need for privacy in large model fine-tuning, particularly with PEFT methods. The integration of large model fine-tuning and differential privacy holds significant promise for future research. However, existing differential privacy techniques, such as DP-SGD [240] and DP-AdamW [241], often result in limited performance and substantial computaitional cost. Therefore, future reasearch should focus on developing methods that preserve privacy while simultaneously optimizing performance and minimizing computational costs. Additionally, exploring scalable, privacy preserving methods tailored to PEFT methods is essential. These advancements will enable secure and efficient fine-tuning of large models, ensuring robust privacy protections.

### v3

@@ V Future Directions
    PEFT methods for multi-objective tasks: Current PEFT methods mainly focus on optimizing for single objectives (e.g., task accuracy), but real-world applications often require balancing multiple objectives (e.g., privacy, fairness, latency). For example, in healthcare, models must preserve patient privacy while maintaining diagnostic accuracy. Existing methods like LoRA or Adapters lack explicit mechanisms to handle such trade-offs. In recent work [263], the authors addressed the program repair task by incorporating a dual-objective optimization framework, wherein the two objectives were combined through linear weighting with manually predefined coefficients to formulate the model’s loss function. Although this study presents a straightforward and effective approach to PEFT for multi-objective tasks, determining the optimal weighting coefficients remains non-trivial. This limitation highlights the need for developing more flexible and task-adaptive methodologies to enhance the robustness and generalizability of such approaches.
@@
    PEFT methods in multimodal learning: Multimodal models (e.g., vision-language models) face unique challenges in aligning heterogeneous data streams (text, images, audio). Current PEFT methods (e.g., adapters) are primarily designed for unimodal LLMs, leading to suboptimal performance in tasks like visual question answering. Recent work on CLIP adaptations [264] highlights the need for modality-specific parameter-efficient tuning to bridge domain gaps. Multimodal learning has emerged as one of the most prominent research topics in contemporary machine learning. However, significant challenges persist in effectively integrating cross-modal information through parameter-efficient fine-tuning (PEFT) approaches, particularly in achieving optimal inter-modal alignment and representation learning while maintaining computational efficiency.
@@
    Automated design of adapter modules: Adapter architectures (e.g., bottleneck layers) rely on manually tuned hyperparameters (e.g., dimension, placement), which limits scalability. Neural Architecture Search (NAS) techniques [265] could automate adapter design, optimizing for both parameter efficiency and task performance. However, the extensive design space of adapter modules significantly compromises the efficiency of NAS approaches. This limitation necessitates further investigation into more efficient and flexible automated design methodologies that can navigate the complex parameter space effectively while maintaining architectural optimality.
@@
    Heuristic search strategies for hybrid PEFT methods: Hybrid methods (e.g., combining LoRA and adapters) often rely on trial-and-error combinations, lacking principled strategies. For example, in paper [197], the authors, under a predefined design space, conduct numerous experiments to determine an ideal hybrid strategy. However, the optimal hybrid strategy may not be included within this artificially predefined design space. Therefore, introducing heuristic search strategies to find the best hybrid strategy is a promising direction for future research.
@@
    Continual learning for PEFT methods: Deployed models must adapt to evolving data distributions (e.g., user preferences in chatbots). Traditional PEFT lacks mechanisms to prevent catastrophic forgetting. Current work [266] proposed a method for task-free online continual learning that dynamically adapts pretrained Vision Transformer models by adding new low-rank adaptation parameters when the loss surface plateaus, indicating data distribution shifts, and uses online weight regularization to mitigate catastrophic forgetting. The experimental results presented in this paper demonstrate significant performance improvements through the application of LoRA, establishing a valuable reference framework for investigating continual learning paradigms in other types of PEFT methodologies.
@@
    Differential privacy for PEFT methods: Different downstream tasks often involve varying levels of sensitve and personal data, which further emphasizes the need for privacy in large language model fine-tuning, particularly with PEFT methods. The integration of large language model fine-tuning and differential privacy holds significant promise for future research. However, existing differential privacy techniques, such as DP-SGD [271] and DP-AdamW [272], often result in limited performance and substantial computaitional cost. Therefore, future reasearch should focus on developing methods that preserve privacy while simultaneously optimizing performance and minimizing computational costs. Additionally, exploring scalable, privacy preserving methods tailored to PEFT methods is essential. These advancements will enable secure and efficient fine-tuning of large language models, ensuring robust privacy protections.